Session: Oct 2008 (session 3)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Sun Oct 19 02:08:31 2008) Success :

Onsala (Michael Lindqvist at Sun Oct 19 08:30:38 2008) Minor failures :
Rain first hours. Some bad scan_checks. Eg. 2008.292.23:27:08.15/scan_check/4,gw019b_on_no0068,mark4,64,2008y292d23h21m2.350s,357.8s,16.000000,73728 A few bad Tsys. Problem with firing of the noisediode. Mark5 error: 2008.293.06:48:32.68:setup01 2008.293.06:48:45.50/mk5/!play_rate = 0 ; 2008.293.06:48:45.50/mk5/!mode = 0 ; 2008.293.06:48:50.89?ERROR m5 -900 Delayed or not mounted 2008.293.06:48:50.89?ERROR m5 -906 MARK5 return code 6: inconsistent or conflicting request 2008.293.06:48:50.97/mk5/!mode? 0 : mark4 : 64 : mark4 : 64 : S : 1 ; 2008.293.06:48:50.97:!2008.293.06:48:30 2008.293.06:48:50.97:preob FM timed out: 2008.293.06:50:00.06:postob 2008.293.06:50:02.30#matcn#debug: #92 /STA[lf]00x00a8 0x0000 0x0a2 40 1 32000 12 2008.293.06:50:04.99#matcn#debug: #92 /SHO 0 0[lf]map[0].stack[1].head[2] = x10 x12 x50 x52 x30 ax32 x70 x72 x11xd31 x33 x71 x73 2008.293.06:50:07.69#matcn#debug: #92 /SHO 0 16[lf]maap[0].stack[1].0 x02 x40 x42 x230 x41 x43 x21 x2 2008.293.06:50:10.42#matcn#debug: #92 /SHO 0 16[lf]maap[0].stack[1].02 x40 x402 x20 x22 x60 x1a x43 x21 x23 x6 2008.293.06:50:12.96#matcn#debug: #92 /SHO 0 16[lf]map[0].staack[1].head[18] 23 x20 x22 x60 x6 ax43 x21 x23 x61 2008.293.06:50:12.96?ERROR ma -4 Device fm timed-out on response.

Medicina (orfei at Sun Oct 19 09:47:31 2008) Success :

Urumqi (Hua Zhang at Mon Oct 20 02:42:35 2008) Failure :
The receiver Lo is wrong for gw019b.skd.In fact,The Lo is 8100MHz,it is not 8080MHz. I am sorry, because I am not find wrong in time.

Yebes (Pablo de Vicente/Rebeca Soria at Mon Oct 20 09:21:30 2008) Minor failures : RECO
Rain during part of the observation Mark5B hung at 01:30 approx. and I had to restart completely the computer because I could not switch off the disk pack. This failure might be related to the fact that I extracted the full disk pack in bank A and inserted a new one (without turning the key) by then. After rebooting I extracted the disk pack with only 90 minutes of data and inserted a new one. At 6:00 the disk bank switch was not automatic and had to be done manually. The disk pack I extracted and with partial information has been tested this morning and is correct.

Effelsberg (Uwe Bach at Mon Oct 20 09:23:09 2008) Success :

Arecibo (T. Ghosh at Tue Oct 21 12:25:26 2008) Success :

Westerbork (Geert Kuper at Wed Oct 22 21:10:53 2008) Minor failures : PHAS FS
Wb used in phased array mode. X/Y phase difference not optimal. We also had to resync the fs time a couple of times, because the fs computer clock was synced by our ntp server but the 'rate model' was still on.

Robledo (Esther Moll at Thu Oct 30 15:12:03 2008) Success :
1. Formatter was on sync. 2. No antenna problems, all sources tracked. 3. SR properly tracking 4. Mark5 recorder#1 performance OK. All scan_check correct. 1932787995096 bytes of data recordedon WSRT-032/2000/1024 module. 1954244354040 bytes of data recordedon WSRT-064/2000/1024 module. 252725108376 bytes of data recordedon WSRT-001/1280/1024 module. 5. All video converters were affected by rain from 02:47UT until EOT. 6. Phase cal tones on. No anomalies observed. 7. Only one noise diode can be configured from EAC; only X-band LCP Tsys monitoring provided. Antabfs file generated. Log and antabfs file transferred to vlbeer. 8. gps-fmout provided at experiment log.

Noto (P. Cassaro at Mon Nov 3 09:25:06 2008) Minor failures : TSYS
THe bbc#8 has been in overflow during a long part or the experiment.